### v1

# FUSION IS ALL YOU NEED : FACE FUSION FOR CUSTOMIZED IDENTITY-PRESERVING IMAGE SYNTHESIS

Abstract
Text-to-image (T2I) models have significantly advanced the development of artificial intelligence, enabling the generation of high-quality images in diverse contexts based on specific text prompts. However, existing T2I-based methods often struggle to accurately reproduce the appearance of individuals from a reference image and to create novel representations of those individuals in various settings. To address this, we leverage the pre-trained UNet from Stable Diffusion to incorporate the target face image directly into the generation process. Our approach diverges from prior methods that depend on fixed encoders or static face embeddings, which often fail to bridge encoding gaps. Instead, we capitalize on UNet’s sophisticated encoding capabilities to process reference images across multiple scales. By innovatively altering the cross-attention layers of the UNet, we effectively fuse individual identities into the generative process. This strategic integration of facial features across various scales not only enhances the robustness and consistency of the generated images but also facilitates efficient multi-reference and multi-identity generation. Our method sets a new benchmark in identity-preserving image generation, delivering state-of-the-art results in similarity metrics while maintaining prompt alignment.
1 Introduction
The synthesis of images based on text prompts has been made possible by the extraordinary capabilities of recently developed large text-to-image (T2I) models. The generated image has high quality and can be in diverse contexts. The strong semantic prior that is acquired from a vast collection of image-caption pairs is one of the primary benefits of these models. However, the output domain of such models is restricted by the bottleneck of the quality and the number of image-caption pairs. The generated contents are hard to control only based on text prompts, and the expressiveness of text is limited, which further constrains the content generalization. Moreover, models that embed the text in an integrated language-vision space are unable to precisely reconstruct the appearance of specific subjects; they can only generate variations of the image content. Existing work like DreamBooth Ruiz et al. (2023) aim to tackle the issue of subject inconsistent generation of T2I models. DreamBooth can be applied to different T2I models, enabling the model to produce more personalized and finely calibrated outputs after training with class-specific prior preservation loss on three to five images of a subject. However, DreamBooth need unique identifier for personalized training with specific subject and is not dedicated for human subject. IP-Adapter Ye et al. (2023) is proposed for mitigating the restricted information expressiveness of text prompt by utilizing the image as the
prompt. It is mainly based on decoupled cross-attention mechanism to separate image feature and text feature and training these newly added layers, the input together The cross attention is coming from encoders (CLIP Radford et al. (2021) or Face Embedding model) and reshaped by a projection model to match target shape if necessary . IP-Adapter can control the model generation for various content generation based on specific image prompt while ensuring the content alignment to the text. While previous methods such as IP-Adapter Ye et al. (2023), InstantID Wang et al. (2024) achieve good image results, however, the generated images are lacking of face fidelity with less controllability. Moreover, they often generate cartoonish looking faces and not completely realistic.
As shown in Figure 1, current ID-preserving methods have less control of consistent identity appearance generation. Moreover, the text information provided by prompt is also limited in previous methods and they cannot be fully reflected in the generated content. For example, with “Cry” as text prompt, the generated images from IPA-FaceID-Plus and InstantID fails to accurately express the ‘crying’ face appearance. They are negatively affected by the appearance of reference image. In the contrast, our method is able to generate accurate facial expression indicated by the text prompt. In this work, we introduce a novel method that aims to address these limitations and capable of generating high fidelity face images and achieving better results in terms of the ID similarity without losing the prompt alignment. Our work utilizes the face image as the reference overcoming drawback of encoders like CLIP Radford et al. (2021) by fusing the reference face image with the one to be generated directly during the cross-attention process. The new keys and values are introduced for learning the face information of reference image, which are combined with text features during training. Such a design enables our method to achieve high image fidelity while remaining efficient for training.
Our main contributions are summarized as follows: • We introduce a new method for enhancing ID preserving ability of existing T2I models by incorporating reference face image directly into diffusion process which achieves SOTA result in terms of ID similarity. Our method shows the superiority in face controllability generation through the corresponding text prompt without degradation on face alignment. • OurOur main contributions are summarized as follows: • We introduce a new method for enhancing ID preserving ability of existing T2I models by incorporating reference face image directly into diffusion process which achieves SOTA result in terms of ID similarity. Our method shows the superiority in face controllability generation through the corresponding text prompt without degradation on face alignment. 2 • Our approach can be also extended for multiple identity generation and supports multiple face references. • We conduct extensive experiments comparing our method with other SOTA identitypreserving techniques regarding facial similarity and text-image alignment, offering empirical validation for the efficacy of the proposed approach. approach can be also extended for multiple identity generation and supports multiple face references. • We conduct extensive experiments comparing our method with other SOTA identitypreserving techniques regarding facial similarity and text-image alignment, offering empirical validation for the efficacy of the proposed approach.
Figure 1: Comparison with other SOTA methods in terms of facial expression generalization and prompt alignment, first row is the input identities, it’s noticed that IPA-FaceId-plus Ye et al. (2023) loses the Face fidelity while InstantId Wang et al. (2024) loses a bit of the prompt alignment, on the other hand our’s method clearly maintains both.
Figure 2: Comparison between using n references to guide the generation by our methods, first column represents the references identities, second row is The guidance using only one image offsets the references, second is using 2 references And third using all 3, We notice that generated face is mix of the features in the references.
2 Related Work
2.1 TEXT-TO-IMAGE DIFFUSION (T2I) MODELS
Recently, diffusion models (DM) have emerged as the current state-of-the-art in text-to-image (T2I) creation. Diffusion-based T2I synthesis has garnered significant attention on social media. A variety of studies on the T2I diffusion model has been published in the past few years, with additional research anticipated in the near future. Diffusion model can be classified into two groups according to the location (e.g., in pixel space and latent space) of diffusion prior. GLIDE Nichol et al. (2021) and Imagen Saharia et al. (2022) generate images directly from the high-dimensional pixel level with pixel space diffusion prior. On the other hand, other research suggests that the diffusion model can be trained on a low-dimensional space after the image has been compressed. Stable diffusion Rombach et al. (2022), VQ-diffusion Gu et al. (2022), and DALL-E 2 Ramesh et al. (2022) are examples of methods that belong to the latent space group. GLIDE adopts classifier free guidance in T2I by replacing the original class label with text, which is preferred by human evaluators than CLIP Radford et al. (2021) guidance. Imagen utilizes pre-trained language models as text encoders for T2I generation and introduces dynamic threshold to classifier-free guidance. Rather than directly training the diffusion model in pixel space, Stable Diffusion optimizes pre-trained autoencoders to encode an image and its text description into vectors that are in close proximity. Cross-attention layers are introduced to train the model in this latent vector space. Stable Diffusion suggests a hyperparameter guidance scale, which a higher guidance scale leads to a greater degree of adherence to the text prompt. VQ-diffusion introduces a non-autoregressive manner to model by leveraging the VQ-VAE to transfer image in the format of discrete image tokens into latent space. DALL-E 2 utilizes the pre-trained CLIP model Radford et al. (2021) that learned the link between textual semantics and visual representations to map the inputs to latent space. Then the modified GLIDE model is applied to stochastically decode CLIP image embeddings from latent space.
2.2 SUBJECT-DRIVEN IMAGE GENERATION
Subject-specific text-to-image generation utilizes a restricted collection of photos of a specific subject to produce tailored visuals derived from textual descriptions. DreamBooth Ruiz et al. (2023) is one of subject-driven generation methods, that is based on fine-tuning T2I models to integrate subject information into various output context domains. DreamBooth constructs a specific prompt with a unique identifier for the pre-defined subject e.g., “A [V ] dog”. Besides, the class-specific prior preservation loss is introduced for maintaining semantic information expressiveness of pre-trained model and enabling diverse subject-specific image generation. ProFusion Zhou et al. (2023) is a regularization-free approach for customized T2I generation. ProFusion includes two parts called PromptNet and Fusion Sampling. PromptNet is used to encode words into the embedding space while conditioned on reference image. Furthermore, fusion sampling is applied to align the generation with the input text context and preserve detailed visual features of the input image. E4T Gal et al. (2023) is an encoder-based solution that uses a word-embedding encoder and weight offsets module for training the personalized T2I models in order to create images with specific concepts from text prompts.
2.3 ID PRESERVING IMAGE GENERATION
Current T2I models have limitations for ID-preserving image generation. The subject of the generated image is not being controlled and prompt engineering cannot promise such purpose. IP-Adapter Ye et al. (2023) utilizes the image prompt to facilitate customized image generation. The input image is encoded into image features by pretrained CLIP modelRadford et al. (2021). Then the image features are further mapped by a linear mapping model before feeding into cross-attention layers. IP-Adapter-FaceID is based on IP-Adapter which is dedicated for realistic face image synthesis by using face ID embedding from face recognition model. For accurate portrait face generation, IPAdapter-FaceID-Portrait is proposed which supports input multiple face images. The generated face embedding are concatenated together as the final input face ID embedding. For more realistic and ID-preserving face generation, IP-Adapter-FaceID-Plus takes directly the reference image as input for diverse face image generation. Overall, all different IP-Adapter methods primarily rely on CLIP image encoder. Compared with the IP-Adapter, InstantID Wang et al. (2024) is proposed for instant personalized image generation which is compatible with community pretrained models without the requirement of fine-tuning. InstantID introduces a trainable lightweight adapted module for projecting face features from the face encoder into the same space of text features. Besides, the IdentityNet is proposed to encode semantic face features from the reference ID image and face embedding with additional input such as landmarks for weak spatial control. Modified ControlNet Zhang et al. (2023) is used as IdentityNet to exploit spatial information of reference face image by only using five facial keypoints and replace text prompts with ID embedding as the condition in cross-attention layers. Arc2Face Papantoniou et al. (2024) is an identity-conditioned foundation model for photorealistic face image synthesis based on only the ID embedding from pretrained ArcFace networks Deng et al. (2019). It is proposed to generate high quality face dataset like Flickr-Faces-HQ Dataset (FFHQ) Karras et al. (2019) but with more identities. Arc2Face uses ID embedding that extracted from a cropped face image after feeding into ArcFace model. Then the ID embedding is mapped into CLIP latent space by applying text encoder and being added into cross-attention layers of finetuned UNet together with text features from other words. The used UNet is especially fine-tuned with high-resolution datasets.
3 Method
3.1 Overview
We propose a novel method for customized, identity-preserving image generation with the following objectives: • Maintain original prompt alignment: Ensure the generated images remain aligned with the original prompts across diverse image generation scenarios. • Generalize across face views and expressions: Enable the model to generalize effectively to different face views and expressions using just a single image sample as input. • Achieve realistic appearance and identity fidelity: Produce images that not only look realistic but also accurately preserve the facial identity of the reference subject. • Efficient training: Design the method requires limited training resources and time. Unlike other approaches that rely on fixed encoders (e.g., IP-Face-Adapter-plus uses a CLIP image encoder as described in Radford et al. (2021)) or face embeddings, our method leverages the visual knowledge embedded in the pre-trained UNet of a Stable Diffusion model Rombach et al. (2022). We input the target face identity image into the UNet, concatenated with the image to be generated. Cross-attention is then applied between these two inputs at various scales and layers within the UNet. This approach facilitates more robust and consistent generation results. To allow the use of face image in guidance, we modify the original cross-attention layers of the UNet. Specifically, we introduce new keys and values layers that take the face image as input and generate new keys and values encoding the face identity. These are then concatenated with the keys and values generated by the text layers, as illustrated in Figure 2. By avoiding the use of fixed encoders and face embeddings, which can suffer from encoding gaps, our method provides a more flexible and effective solution for identity-preserving image generation. The integration of the target face identity directly into the UNet’s cross-attention mechanism ensures that the generated images retain the desired facial characteristics while maintaining alignment with the input prompts and generalizing well across different face views and expressions.
3.2 FUSED ATTENTION
The cross-attention mechanism is applied at various layers of the UNet to integrate the input condition prompt c with the hidden state h, which represents the latent vector of the image being generated at a specific layer within the UNet. The queries $Q_{h}$, keys $K_{c}$, and values $V_{c}$ are computed as follows:
$Q_{h}=Linear_{q}(h),K_{c}=Linear_{k}(c),V_{c}=Linear_{v}(c)$ (1)
Here, $K_{c}$ and $V_{c}$ are of dimension $R^{N_{c}\times D}$, and $Q_{h}$ is of dimension $R^{N_{h}\times D}$. The functions Linearq, Lineark, and Linearv represent the original linear layers from the Stable Diffusion (SD) model Rombach et al. (2022). The keys and values derived from the text prompt are used to compute the attention score, which is subsequently multiplied with the values. This mechanism ensures that the guidance provided by the text is effectively integrated into the generation process. To enhance the identity preservation and generalization capabilities, we modify these linear layers to incorporate the target face identity. Specifically, we introduce additional layers that process the face identity image and generate new keys and values that encode the facial features. These new keys and values are then concatenated with the original keys and values derived from the text prompt. We append two additional layers, $Linear_{k_{f}}$ and $Linear_{v_{f}}$, to the attention mechanism. These layers represent the only trainable components of our proposed pipeline. For a face image latent hidden state hf at a specific attention layer within the UNet, new keys and values that encapsulate the face identity are computed as follows:
$K_{f}=Linear_{k_{f}}(h_{f}),V_{f}=Linear_{v_{f}}(h_{f})$ (2)
Here, $K_{f}$ and $V_{f}$ are of dimension $R_{Nh\texttimes D}$. The final keys K‘ and values V ‘ , which incorporate both the prompt and the face identity, are obtained by concatenating the original keys and values with the newly computed face identity keys and values:
$K`=[K_{c};K_{f}],V`=[V_{c};V_{f}]$ (3)
In this equation, K‘ and V‘ are of dimension $R^{(Nc+Nh)\times D}$. By integrating the face identity through these additional layers and concatenation, our method ensures that the generated images not only align with the input text prompts but also accurately preserve the facial characteristics of the target identity. This approach allows for more precise control over the generated images, enabling the creation of high-quality, identity-preserving results. The efficiency of our method is maintained by limiting the trainable components to just the two additional layers, Linearkf and Linearvf . This design choice reduces the computational overhead and training time required, making our approach practical for real-world applications. The attention scores are computed as follows:
$S=Softmax(M+\alpha\cdot Q`K`^{T})$ (4)
Here, M denotes the attention mask, $\alpha$ is a scale term, and S represents the attention scores that incorporate both the text and the face image. The attention scores matrix S has dimensions $R^{Nh\times(N_{c}+N_{h})}$ Finally, the output hidden state h is obtained by multiplying the attention scores with the concatenated values:
$h=SV$ (5)
The resulting hidden state h has dimensions $R^{N_{h}\times D}$. As mentioned in Section 2, previous methods often separate the identity (ID) guidance into a separate attention layer, where the result of the ID attention is added to the output of the original crossattention layer guided by the text. In contrast, we find that our approach of concatenating the keys and values from the face identity directly into the original cross-attention layer provides stronger and more consistent ID guidance. This method avoids potential conflicts that can arise from combining two separate guiding attention layers. Additionally, our approach eliminates the need to project the face image to match a specific target shape. Instead, the concatenated keys and values are seamlessly integrated during the multiplication process, and the hidden state is restored to its original size without any additional projections. By concatenating the keys and values, our method ensures that the identity information is effectively integrated into the attention mechanism, leading to more accurate and consistent preservation of facial characteristics in the generated images. simplifies the architecture but also improves the overall performance and efficiency of the model.
3.3 MULTIPLE REFERENCE
Overall, we believe that a single image clearly showing the identity instance is sufficient for identitypreserving generation. However, there are cases where using multiple references can yield better results, especially when the face in the adapter is occluded or different features are displayed across different reference images. Our method can easily accommodate the use of multiple references. For identities $f1,\dots,fn$ with corresponding hidden states $h1,\dots,hn$, the keys and values can be obtained using Equation 2. Equation 3 is then modified as follows:
$K`=[K_{c};K_{f1};\dots;K_{fn}],V`=[V_{c};V_{f1};\dots;V_{fn}]$ (6)
This modification allows for the integration of any number of identities without the need for further training or fine-tuning. We showcase sample results in Figure 3, demonstrating the effectiveness of using three identities for improved generation. The same principle can be applied to generate an image with an identity that is a mixture of distinct identities, a process commonly known as face morphing. More details on this application can be found in the appendix. By extending our method to handle multiple references, we enhance the flexibility and robustness of the identity-preserving generation process. This capability allows the model to leverage information from multiple sources, leading to better handling of occlusions and variations in facial features across different images.
3.4 MULTIPLE IDENTITIES
Handling multiple identities in image generation remains an open challenge. In this work, we propose a novel solution, that combines our previously mentioned method to address this issue. The goal is to generate an image containing (n) distinct identities ${f1,\dots,fn}$ based on input references and a given prompt. Given masks ${m1,\dots,mn}$ where mi indicates The location of The face for identity i, We scale each mask and feed to each cross-attention layer In The UNet, The scaling issue done based one the scale of the hidden state input To the attention and for some layer with scales si We have masks ${m1si,\dots,mnsi}$. For each of the face identities We can get get keys And values And combined them similar to Equation 6, if We applied attention like this we will obtain faces that are mix off all identities features (i.e. face morphing), we instead use a customized attention mask M, where The mask can is defined as:
$M_{q,k}=\begin{cases}-\inf&\text{if }isFaceKey(k)\times identity(k)=i\wedge\exists m_{j}:m_{j_{q}}=true\wedge j\neq i\\ 1&\text{otherwise }\end{cases}$ (7)
where q indicates the query and k is the key. This approach ensures that all identities are present in the generated image. We demonstrate the effectiveness of our method with graphical results showing example input identities in Figure 4. These results highlight the success of our approach in generating images conditioned on multiple face identities.
Figure 3: The overall pipeline of our proposed method. We first concatenate the target face image with the image to be generated, We modify the cross attention layers to include new keys any values layers their output is concatenated with the output of the original ones from text prompt.
4 Experiments
Figure 4: Multiple face identities image generation using our proposed method, first row represents the three input distinct identities, following row contains sample results for different prompts, depthmap based ControlNet Zhang et al. (2023) was used for the pose.
Figure 5: Comparison with other SOTA methods, with prompt used as ”astronaut in a garden” and the first row represents the input target faces identities, the following rows are results from IP-faceAdapter Ye et al. (2023), IP-face-Adapter-plus Ye et al. (2023) and ours respectively. Identities are selected from ffhq datasetKarras et al. (2019)
4.1 EXPERIMENT SETTINGS
In our work, we utilize Stable Diffusion Rombach et al. (2022) as the base model, specifically the Realistic Vision v4.0 checkpoint1 , we train a subset from the LAION-face dataset Zheng et al. (2022) consisting of approximately 80000 text-image pairs. Our training and all the experiments are done on 2 Tesla V100 GPUs (30 GB) with batch size of 1 per GPU. For the other models compared in our work, we use InstantID Wang et al. (2024) which uses SDXL model (offering an advantage over our method), as for IPA-FaceId, we use IPA-FaceID-plusv2 Ye et al. (2023) and both using Realistic Vision v4.0 checkpoint as the base model, for judging the CLIP score, we generate listing of 100 identity conditioned on different prompts using LLMs. We provide two versions of our method, the second version is trained for more timestep and larger dataset.
4.2 Evaluation Metrics
Previous research on identity-preserving imagine generation tasks Wang et al. (2024); Ye et al. (2023); Papantoniou et al. (2024) lacks a defined metric for assessing identity preservation. We employ cosine distance to assess the ID-conditioning capability of various approaches in our study. Face embeddings are acquired by feeding generated images into pre-trained face recognition models. We compare the face embedding of the reference image with those from the created image; a reduced distance signifies superior ID-consistent generating performance. Additionally, we calculate the CLIP score using the CLIP modelRadford et al. (2021) to assess content alignment in face image generation across several scenarios with varied text prompts. The higher CLIP score indicates high  [^1] https://huggingface.co/SG161222 text-image alignment. For the evaluation of image quality, we employ SSIM and PSNR to compare the generated images with the reference images. In order to measure the face fidelity, we focus on calculate those metrics based on the face region only by cropping the faces from generated images.
Figure 6: Comparison with other SOTA methods, the first column are the input identities along with input styles, From left to right, we show the results of IPA Ye et al. (2023), IPA-FaceID-SDXL, IPAFaceID-Plus, InstantId and our’s, it’s noticed that previous methods unlike our’s can suffer from lack of Face fidelity or style degradation.
4.3 QUALITATIVE RESULTS
In this section, we mainly evaluate our proposed method for visual comparison in terms of ID preservation. As showns in Figure 5, we generate images based on multiple identity face images with the same text prompt ”astronaut in a garden”. IPA-FaceID tend to generate face images with low ID preservation and it sometimes fails to maintain the gender of the reference face image. On the other hand, IPA-FaceID-Plus has better performance on maintaining the gender, age of input identity, however, it still loses some original facial attributes such as facial emotion and hairstyles. In the contrast, our method generates face images have high level of ID preservation (preserving facial expression, gender, age and detailed feature such as hairstyles and skincolor). Specifically, in order to investigate the accurate facial expression generation, we generate images with different text prompts like ”Anger”, ”Cry”, ”Laugh” to control the face emotion of generation. As depicted in Figure 1, InstantID generates images with less emotion alignment with associated text prompt while IPA-FaceID-Plus can generate intend emotion. However, IPA-FaceID-Plus has less control of ID preservation with exaggerated face expression in random styles (noticeable changes in facial structures and diverse backgrounds). Our method provides the strong control of facial expression based on corresponding text prompt. Especially in the case of ”Cry” prompt, even though the reference image already with ‘smile’ or ‘laugh’ facial expression, our method still can address the importance of text prompt to the final image generation. Our method produces ID-consistent face images while has more natural-looking face expression representations without suppressing the contribution of text prompt to the generated images. In Figure 6, we aim to compare the ID preservation in different image styles generation. Same as previous work Wang et al. (2024), we use a set of text prompts to indicate various image styles like ”Line art”, ”Faded Film” and ”Watercolor”. Our method is able to generate comparable results with other SOTA methods while maintaining high degree of face and style alignment together.
4.4 QUANTITATIVE RESULTS
Besides visual comparisons, we also numerically evaluate our method based on CLIP score, SSIM, PSNR and cosine distance using three models namely VGG-Face, Facenet512 and ArcFace. We utilize a subset of FFHQ Karras et al. (2019) dataset high quality face images as the input identities in these experiments.
Table 1 shows the numerical results of cosine distance on different methods. Firstly, in terms of the face alignment, it’s clear that our methods achieve significantly higher alignment than the other methods on all face recognition models. The faces in our generated image maintain high level of face features from original reference identity and can be easily used for accessing different face recognition systems. In Table 2, we judge The face quality And prompt alignment offering different methods, IPA-FaceId has the highest CLIP score, however it’s ID similarity is very low compared together ours and other methods, Ours v1 And IPA-FaceID-Plus maintains almost similar CLIP score with IPA-FaceID-Plus being slightly better, Ours v2 is worse In terms of CLIP but it achieves significantly higher identifier similarity. We exclude InstantId Wang et al. (2024) from CLip Score due to the fact that it requires pose conditioning And the base model for the public checkpoint issue SDXL based, which will make The comparison not plausible. We also compare The face quality using PSNR and SSIM metrics, our method achieves benefits result on both further demonstrating that our method is capable of generating high quality realistic faces with high ID fidelity.
5 LIMITATIONS AND FUTURE WORK
We have observed that fine facial features tend to degrade when the face occupies a small portion of the overall image. This issue primarily stems from the limitations of Stable Diffusion Rombach et al. (2022). Additionally, when our method is used in conjunction with ControlNet Zhang et al. (2023), it results in reduced identity fidelity, which can begin resolved by finetuning. In our future work, we aim to address these challenges by enhancing the ability to maintain fine details and improving control in scenarios involving the generation of multiple identities.
6 CONCLUSION
In this paper, we present a novel method for ID-preserving image generation that excels in maintaining high identity fidelity while producing realistic facial images. Our approach demonstrates robust generalization across various facial views and expressions. Additionally, we propose innovative solutions for handling multiple references and multi-identity generation, effectively bridging encoder domain gaps and leveraging multi-scale guidance for enhanced performance. Our method achieves new state-of-the-art results in preserving identity similarity, as evidenced by our extensive experimental evaluations. Furthermore, it maintains good alignment with given prompts, setting a new benchmark in the field of identity-preserving image generation.
Our proposed methods is capable of high prompt alignment with multi-views Face deduction from one sample input, while preserving the target identity with high fidelity 2, we can clearly see that other methods can fail in identity preservation on certain difficult reference inputs and loses prompt alignment, generating same face pose And landmarks is also another issue, we can see that our proposed pipeline and method of training overcome these issue as shown in 5.

### v2

@@
 1 Introduction
-The synthesis of images based on text prompts has been made possible by the extraordinary capabilities of recently developed large text-to-image (T2I) models.
-The generated image has high quality and can be in diverse contexts.
-The strong semantic prior that is acquired from a vast collection of image-caption pairs is one of the primary benefits of these models.
-However, the output domain of such models is restricted by the bottleneck of the quality and the number of image-caption pairs.
-The generated contents are hard to control only based on text prompts, and the expressiveness of text is limited, which further constrains the content generalization.
+The synthesis of images based on text prompts has been made possible by the extraordinary capabilities of recently developed large text-to-image (T2I) models, the generated image has high quality with diverse contexts, the strong semantic prior that is acquired from a vast collection of image-caption pairs is one of the primary benefits of these models.
+However, the output domain of such models is restricted by the bottleneck of the quality and the number of image-caption pairs, the generated contents are hard to control only based on text prompts, and the expressiveness of text is limited, which further constrains the content generalization.
 Moreover, models that embed the text in an integrated language-vision space are unable to precisely reconstruct the appearance of specific subjects; they can only generate variations of the image content.
-Existing work like DreamBooth Ruiz et al.
-(2023) aim to tackle the issue of subject inconsistent generation of T2I models.
-DreamBooth can be applied to different T2I models, enabling the model to produce more personalized and finely calibrated outputs after training with class-specific prior preservation loss on three to five images of a subject.
+Existing work like DreamBooth (Ruiz et al., 2023) aim to tackle the issue of subject inconsistent generation of T2I models, DreamBooth can be applied to different T2I models, enabling the model to produce more personalized and finely calibrated outputs after training with class-specific prior preservation loss on three to five images of a subject.
 However, DreamBooth need unique identifier for personalized training with specific subject and is not dedicated for human subject.
-IP-Adapter Ye et al.
-(2023) is proposed for mitigating the restricted information expressiveness of text prompt by utilizing the image as the
-prompt.
-It is mainly based on decoupled cross-attention mechanism to separate image feature and text feature and training these newly added layers, the input together The cross attention is coming from encoders (CLIP Radford et al.
-(2021) or Face Embedding model) and reshaped by a projection model to match target shape if necessary .
-IP-Adapter can control the model generation for various content generation based on specific image prompt while ensuring the content alignment to the text.
-While previous methods such as IP-Adapter Ye et al.
-(2023), InstantID Wang et al.
-(2024) achieve good image results, however, the generated images are lacking of face fidelity with less controllability.
+IP-Adapter (Ye et al., 2023) is proposed for mitigating the restricted information expressiveness of text prompt by utilizing the image as the prompt, it is mainly based on decoupled cross-attention mechanism to separate image feature and text feature and training these newly added layers, the input to the cross attention is coming from encoders (CLIP (Radford et al., 2021) or Face embedding model) and reshaped by a projection model to match target shape if necessary.
+IP-Adapter can control the model generation for various content generation based on specific image prompt while ensuring the content alignment to the text, InstantID (Wang et al., 2024) is another method newly proposed for face-identity conditioned image generation also utilizing face embeddings, while previous methods such as IP-Adapter (Ye et al., 2023) and InstantID (Wang et al., 2024) achieve good image results, however, the generated images are lacking in terms of face fidelity with less controllability.
 Moreover, they often generate cartoonish looking faces and not completely realistic.
-As shown in Figure 1, current ID-preserving methods have less control of consistent identity appearance generation.
-Moreover, the text information provided by prompt is also limited in previous methods and they cannot be fully reflected in the generated content.
-For example, with “Cry” as text prompt, the generated images from IPA-FaceID-Plus and InstantID fails to accurately express the ‘crying’ face appearance.
-They are negatively affected by the appearance of reference image.
-In the contrast, our method is able to generate accurate facial expression indicated by the text prompt.
-In this work, we introduce a novel method that aims to address these limitations and capable of generating high fidelity face images and achieving better results in terms of the ID similarity without losing the prompt alignment.
-Our work utilizes the face image as the reference overcoming drawback of encoders like CLIP Radford et al.
-(2021) by fusing the reference face image with the one to be generated directly during the cross-attention process.
-The new keys and values are introduced for learning the face information of reference image, which are combined with text features during training.
-Such a design enables our method to achieve high image fidelity while remaining efficient for training.
-Our main contributions are summarized as follows: • We introduce a new method for enhancing ID preserving ability of existing T2I models by incorporating reference face image directly into diffusion process which achieves SOTA result in terms of ID similarity.
-Our method shows the superiority in face controllability generation through the corresponding text prompt without degradation on face alignment.
-• OurOur main contributions are summarized as follows: • We introduce a new method for enhancing ID preserving ability of existing T2I models by incorporating reference face image directly into diffusion process which achieves SOTA result in terms of ID similarity.
-Our method shows the superiority in face controllability generation through the corresponding text prompt without degradation on face alignment.
-2 • Our approach can be also extended for multiple identity generation and supports multiple face references.
-• We conduct extensive experiments comparing our method with other SOTA identitypreserving techniques regarding facial similarity and text-image alignment, offering empirical validation for the efficacy of the proposed approach.
-approach can be also extended for multiple identity generation and supports multiple face references.
-• We conduct extensive experiments comparing our method with other SOTA identitypreserving techniques regarding facial similarity and text-image alignment, offering empirical validation for the efficacy of the proposed approach.
-Figure 1: Comparison with other SOTA methods in terms of facial expression generalization and prompt alignment, first row is the input identities, it’s noticed that IPA-FaceId-plus Ye et al.
-(2023) loses the Face fidelity while InstantId Wang et al.
-(2024) loses a bit of the prompt alignment, on the other hand our’s method clearly maintains both.
-Figure 2: Comparison between using n references to guide the generation by our methods, first column represents the references identities, second row is The guidance using only one image offsets the references, second is using 2 references And third using all 3, We notice that generated face is mix of the features in the references.
+As illustrated in Figure 1, existing ID-preserving methods struggle to maintain consistent identity appearance across generated images.
+Additionally, the textual information provided in prompts is often poorly reflected in the outputs of previous methods, resulting in incomplete alignment between the prompt and the generated content.
+For instance, when using “Cry” style as the text prompt, models like IPA-FaceID-Plus (Ye et al., 2023) fails in terms of face fidelity and InstantID (Wang et al., 2024) fails to accurately convey the appearance of a crying face and its’ results are often negatively influenced by the reference image’s appearance.
+In contrast, our method addresses these limitations by generating facial expressions that accurately align with the given text prompt.
+Specifically, we propose a novel approach that not only ensures high-fidelity face generation but also enhances identity similarity without sacrificing prompt alignment.
+Our method leverages the reference face image directly, overcoming the limitations of encoders like CLIP (Radford et al., 2021) by integrating the reference image with the generated one during the cross-attention process.
+New keys and values are introduced to capture the reference image’s facial information, which is then fused with text features.
+This design enables our method to produce high-quality, realistic images while remaining computationally efficient for training.
+Our main contributions are summarized as follows:
+  • $\bullet$
+    We introduce a new method for enhancing ID preserving ability of existing T2I models by incorporating reference face image directly into diffusion process which achieves SOTA result in terms of ID similarity, our method shows the superiority in face controllability generation through the corresponding text prompt without degradation on face alignment.
+  • $\bullet$
+    Our approach can also be extended for multiple identity generation and supports multiple face references.
+  • $\bullet$
+    We conduct extensive experiments comparing our method with other SOTA identity-preserving techniques regarding facial similarity and text-image alignment, offering empirical validation for the efficacy of the proposed approach.
+Figure 1: Comparison with other SOTA methods in terms of facial expression controllability and prompt alignment, first row are the input identities.
+It’s worth to notice that IPA-FaceId-Plus (Ye et al., 2023) loses the face fidelity while InstantID (Wang et al., 2024) loses a bit of the prompt alignment.
+On the other hand our method maintains ID preservation with accurate facial expression generation.
 2 Related Work
 2.1 TEXT-TO-IMAGE DIFFUSION (T2I) MODELS
-Recently, diffusion models (DM) have emerged as the current state-of-the-art in text-to-image (T2I) creation.
-Diffusion-based T2I synthesis has garnered significant attention on social media.
-A variety of studies on the T2I diffusion model has been published in the past few years, with additional research anticipated in the near future.
-Diffusion model can be classified into two groups according to the location (e.g., in pixel space and latent space) of diffusion prior.
-GLIDE Nichol et al.
-(2021) and Imagen Saharia et al.
-(2022) generate images directly from the high-dimensional pixel level with pixel space diffusion prior.
-On the other hand, other research suggests that the diffusion model can be trained on a low-dimensional space after the image has been compressed.
-Stable diffusion Rombach et al.
-(2022), VQ-diffusion Gu et al.
-(2022), and DALL-E 2 Ramesh et al.
-(2022) are examples of methods that belong to the latent space group.
-GLIDE adopts classifier free guidance in T2I by replacing the original class label with text, which is preferred by human evaluators than CLIP Radford et al.
-(2021) guidance.
-Imagen utilizes pre-trained language models as text encoders for T2I generation and introduces dynamic threshold to classifier-free guidance.
-Rather than directly training the diffusion model in pixel space, Stable Diffusion optimizes pre-trained autoencoders to encode an image and its text description into vectors that are in close proximity.
-Cross-attention layers are introduced to train the model in this latent vector space.
-Stable Diffusion suggests a hyperparameter guidance scale, which a higher guidance scale leads to a greater degree of adherence to the text prompt.
-VQ-diffusion introduces a non-autoregressive manner to model by leveraging the VQ-VAE to transfer image in the format of discrete image tokens into latent space.
-DALL-E 2 utilizes the pre-trained CLIP model Radford et al.
-(2021) that learned the link between textual semantics and visual representations to map the inputs to latent space.
-Then the modified GLIDE model is applied to stochastically decode CLIP image embeddings from latent space.
+In recent years, diffusion models (DM) have become the state-of-the-art in text-to-image (T2I) generation, attracting significant attention on social media.
+Numerous studies on T2I diffusion models have been published, with continued research expected in the near future.
+Diffusion models can be broadly categorized into two groups based on their operational space: pixel space or latent space.
+Pixel-space models, such as GLIDE (Nichol et al., 2021) and Imagen (Saharia et al., 2022), generate images directly from high-dimensional pixel data using pixel-space diffusion priors.
+In contrast, latent-space models, such as Stable Diffusion (Rombach et al., 2022), VQ-diffusion (Gu et al., 2022), and DALL-E 2 (Ramesh et al., 2022), operate in a compressed, low-dimensional space after the image has been encoded.
+GLIDE introduced classifier-free guidance for T2I, replacing the original class labels with text prompts, which human evaluators found more effective than CLIP-guided models (Radford et al., 2021).
+Imagen builds on this by using pre-trained language models as text encoders and introducing dynamic thresholding for classifier-free guidance.
+Stable Diffusion takes a different approach by optimizing pre-trained autoencoders to map images and their text descriptions into vectors in latent space.
+Cross-attention layers are employed to guide the generation in this compressed space.
+Similar to GLIDE (Nichol et al., 2021), it follows a classifier-free guidance approach and uses a guidance scale hyperparameter, allowing for greater alignment with the text prompt as the scale increases.
 2.2 SUBJECT-DRIVEN IMAGE GENERATION
-Subject-specific text-to-image generation utilizes a restricted collection of photos of a specific subject to produce tailored visuals derived from textual descriptions.
-DreamBooth Ruiz et al.
-(2023) is one of subject-driven generation methods, that is based on fine-tuning T2I models to integrate subject information into various output context domains.
+Subject-specific text-to-image generation utilizes a collection of photos of a specific subject to produce tailored visuals derived from textual descriptions, DreamBooth (Ruiz et al., 2023) is one of subject-driven generation methods, that is based on fine-tuning T2I models to integrate subject information into various output context domains.
 DreamBooth constructs a specific prompt with a unique identifier for the pre-defined subject e.g., “A [V ] dog”.
 Besides, the class-specific prior preservation loss is introduced for maintaining semantic information expressiveness of pre-trained model and enabling diverse subject-specific image generation.
-ProFusion Zhou et al.
-(2023) is a regularization-free approach for customized T2I generation.
-ProFusion includes two parts called PromptNet and Fusion Sampling.
-PromptNet is used to encode words into the embedding space while conditioned on reference image.
-Furthermore, fusion sampling is applied to align the generation with the input text context and preserve detailed visual features of the input image.
-E4T Gal et al.
-(2023) is an encoder-based solution that uses a word-embedding encoder and weight offsets module for training the personalized T2I models in order to create images with specific concepts from text prompts.
+ProFusion (Zhou et al., 2023) and E4T (Gal et al., 2023) adopt a similar approach, fine-tuning a dedicated identity token for each individual.
+However, this necessitates separate fine-tuning for every identity, making them impractical in many scenarios.
 2.3 ID PRESERVING IMAGE GENERATION
-Current T2I models have limitations for ID-preserving image generation.
-The subject of the generated image is not being controlled and prompt engineering cannot promise such purpose.
-IP-Adapter Ye et al.
-(2023) utilizes the image prompt to facilitate customized image generation.
-The input image is encoded into image features by pretrained CLIP modelRadford et al.
-(2021).
-Then the image features are further mapped by a linear mapping model before feeding into cross-attention layers.
+Current T2I models have limitations for ID-preserving image generation, the subject of the generated image is not being controlled and prompt engineering cannot promise such purpose.
+IP-Adapter (Ye et al., 2023) utilizes the image prompt to facilitate customized image generation.
+The input image is encoded into image features by pre-trained CLIP model (Radford et al., 2021), then the image features are further mapped by a linear mapping model before feeding into cross-attention layers.
 IP-Adapter-FaceID is based on IP-Adapter which is dedicated for realistic face image synthesis by using face ID embedding from face recognition model.
-For accurate portrait face generation, IPAdapter-FaceID-Portrait is proposed which supports input multiple face images.
-The generated face embedding are concatenated together as the final input face ID embedding.
-For more realistic and ID-preserving face generation, IP-Adapter-FaceID-Plus takes directly the reference image as input for diverse face image generation.
+For accurate portrait face generation, IP-Adapter-FaceID-Portrait is another variant which supports input multiple face images.
+For more realistic and ID-preserving face generation, IP-Adapter-FaceID-Plus takes directly the reference image along with face embeddings as input for enhanced face fidelity generation.
 Overall, all different IP-Adapter methods primarily rely on CLIP image encoder.
-Compared with the IP-Adapter, InstantID Wang et al.
-(2024) is proposed for instant personalized image generation which is compatible with community pretrained models without the requirement of fine-tuning.
-InstantID introduces a trainable lightweight adapted module for projecting face features from the face encoder into the same space of text features.
-Besides, the IdentityNet is proposed to encode semantic face features from the reference ID image and face embedding with additional input such as landmarks for weak spatial control.
-Modified ControlNet Zhang et al.
-(2023) is used as IdentityNet to exploit spatial information of reference face image by only using five facial keypoints and replace text prompts with ID embedding as the condition in cross-attention layers.
-Arc2Face Papantoniou et al.
-(2024) is an identity-conditioned foundation model for photorealistic face image synthesis based on only the ID embedding from pretrained ArcFace networks Deng et al.
-(2019).
-It is proposed to generate high quality face dataset like Flickr-Faces-HQ Dataset (FFHQ) Karras et al.
-(2019) but with more identities.
-Arc2Face uses ID embedding that extracted from a cropped face image after feeding into ArcFace model.
-Then the ID embedding is mapped into CLIP latent space by applying text encoder and being added into cross-attention layers of finetuned UNet together with text features from other words.
-The used UNet is especially fine-tuned with high-resolution datasets.
+Compared with the IP-Adapter, InstantID (Wang et al., 2024) is proposed for instant personalized image generation and is compatible with community-pre-trained models.
+Similar to IP-Adapter-FaceID-Plus, it relies on face embeddings with decoupled cross-attention layers, allowing the use of visual prompts.
+Additionally, it introduces a module named IdentityNet, responsible for encoding detailed features from the reference face input image with spatial control using ControlNet (Zhang et al., 2023).
+Arc2Face (Papantoniou et al., 2024) is an identity-conditioned foundation model for photorealistic face image synthesis based on only the ID embedding from pre-trained ArcFace networks (Deng et al., 2019).
+It is proposed to generate high quality face dataset like Flickr-Faces-HQ Dataset (FFHQ) (Karras et al., 2019) but with more identities.
+Arc2Face uses ID embedding that extracted from a cropped face image after feeding into ArcFace model, then the ID embedding is mapped into CLIP latent space by applying text encoder and being added into cross-attention layers of fine-tuned UNet together with text features from other words, However its’ primarily goal is to extract facialid features from face embeddings and not suitable for customized ID-preserving image generation.
 3 Method
+Figure 2: The overall pipeline of the proposed method.
+We first concatenate the target face image with the image to be generated for applying attention at different scales, we modify cross-attention layers to include new keys and values layers, their output is concatenated with the output of the original ones from text prompt.
 3.1 Overview
-We propose a novel method for customized, identity-preserving image generation with the following objectives: • Maintain original prompt alignment: Ensure the generated images remain aligned with the original prompts across diverse image generation scenarios.
-• Generalize across face views and expressions: Enable the model to generalize effectively to different face views and expressions using just a single image sample as input.
+We propose a novel method for customized, identity-preserving image generation with the following objectives:
+  • $\bullet$
+    Maintain original prompt alignment: Ensure the generated images remain aligned with the original prompts across diverse image generation scenarios.
+  • $\bullet$
+    Generalize facial features: Enable the model to generalize effectively to different face views and expressions using just a single image sample as input.
+  • $\bullet$
     Achieve realistic appearance and identity fidelity: Produce images that not only look realistic but also accurately preserve the facial identity of the reference subject.
+  • $\bullet$
     Efficient training: Design the method requires limited training resources and time.
-Unlike other approaches that rely on fixed encoders (e.g., IP-Face-Adapter-plus uses a CLIP image encoder as described in Radford et al.
-(2021)) or face embeddings, our method leverages the visual knowledge embedded in the pre-trained UNet of a Stable Diffusion model Rombach et al.
-(2022).
+Unlike other approaches that rely on fixed encoders (e.g., IP-Face-Adapter-plus uses a CLIP image encoder as described in (Radford et al., 2021)) or face embeddings, our method leverages the visual knowledge embedded in the pre-trained UNet of a Stable Diffusion model (Rombach et al., 2022).
 We input the target face identity image into the UNet, concatenated with the image to be generated.
@@
 To allow the use of face image in guidance, we modify the original cross-attention layers of the UNet.
-Specifically, we introduce new keys and values layers that take the face image as input and generate new keys and values encoding the face identity.
-These are then concatenated with the keys and values generated by the text layers, as illustrated in Figure 2.
+Specifically, we introduce new keys and values layers that take the face image as input and generate new keys and values encoding the face identity, these are then concatenated with the keys and values generated by the text layers, as illustrated in Figure 2.
 By avoiding the use of fixed encoders and face embeddings, which can suffer from encoding gaps, our method provides a more flexible and effective solution for identity-preserving image generation.
@@
 Here, $K_{c}$ and $V_{c}$ are of dimension $R^{N_{c}\times D}$, and $Q_{h}$ is of dimension $R^{N_{h}\times D}$.
-The functions Linearq, Lineark, and Linearv represent the original linear layers from the Stable Diffusion (SD) model Rombach et al.
-(2022).
+The functions $Linear_{q}$, $Linear_{k}$, and $Linear_{v}$ represent the original linear layers from the Stable Diffusion (SD) model (Rombach et al., 2022).
 The keys and values derived from the text prompt are used to compute the attention score, which is subsequently multiplied with the values.
 This mechanism ensures that the guidance provided by the text is effectively integrated into the generation process.
-To enhance the identity preservation and generalization capabilities, we modify these linear layers to incorporate the target face identity.
+To facilitate the identity preservation and generalization capabilities, we modify this attention mechanism to incorporate the target face identity.
 Specifically, we introduce additional layers that process the face identity image and generate new keys and values that encode the facial features.
@@
 These layers represent the only trainable components of our proposed pipeline.
-For a face image latent hidden state hf at a specific attention layer within the UNet, new keys and values that encapsulate the face identity are computed as follows:
+For a face image latent hidden state $h_{f}$ at a specific attention layer within the UNet, new keys and values that encapsulate the face identity are computed as follows:
 $K_{f}=Linear_{k_{f}}(h_{f}),V_{f}=Linear_{v_{f}}(h_{f})$ (2)
-Here, $K_{f}$ and $V_{f}$ are of dimension $R_{Nh\texttimes D}$.
-The final keys K‘ and values V ‘ , which incorporate both the prompt and the face identity, are obtained by concatenating the original keys and values with the newly computed face identity keys and values:
-$K`=[K_{c};K_{f}],V`=[V_{c};V_{f}]$ (3)
-In this equation, K‘ and V‘ are of dimension $R^{(Nc+Nh)\times D}$.
+Here, $K_{f}$ and $V_{f}$ are of dimension $R^{N_{h}\times D}$.
+The final keys $K^{{}^{\prime}}$ and values $V^{{}^{\prime}}$ , which incorporate both the prompt and the face identity, are obtained by concatenating the original keys and values with the newly computed face identity keys and values:
+$K^{{}^{\prime}}=[K_{c};K_{f}],V^{{}^{\prime}}=[V_{c};V_{f}]$ (3)
+In this equation, $K^{{}^{\prime}}$ and $V^{{}^{\prime}}$ are of dimension $R^{(Nc+Nh)\times D}$.
 By integrating the face identity through these additional layers and concatenation, our method ensures that the generated images not only align with the input text prompts but also accurately preserve the facial characteristics of the target identity.
 This approach allows for more precise control over the generated images, enabling the creation of high-quality, identity-preserving results.
-The efficiency of our method is maintained by limiting the trainable components to just the two additional layers, Linearkf and Linearvf .
-This design choice reduces the computational overhead and training time required, making our approach practical for real-world applications.
+The efficiency of our method is maintained by limiting the trainable components to just the two additional layers, $Linear_{k_{f}}$ and $Linear_{v_{f}}$ This design choice reduces the computational overhead and training time required, making our approach practical for real-world applications.
 The attention scores are computed as follows:
-$S=Softmax(M+\alpha\cdot Q`K`^{T})$ (4)
+$S=Softmax(M+\alpha\cdot Q^{{}^{\prime}}{K^{{}^{\prime}}}^{T})$ (4)
 Here, M denotes the attention mask, $\alpha$ is a scale term, and S represents the attention scores that incorporate both the text and the face image.
-The attention scores matrix S has dimensions $R^{Nh\times(N_{c}+N_{h})}$ Finally, the output hidden state h is obtained by multiplying the attention scores with the concatenated values:
-$h=SV$ (5)
+The attention scores matrix S has the dimension $R^{Nh\times(N_{c}+N_{h})}$ Finally, the output hidden state h is obtained by multiplying the attention scores with the concatenated values:
+$h=SV^{{}^{\prime}}$ (5)
 The resulting hidden state h has dimensions $R^{N_{h}\times D}$.
-As mentioned in Section 2, previous methods often separate the identity (ID) guidance into a separate attention layer, where the result of the ID attention is added to the output of the original crossattention layer guided by the text.
+As mentioned in Section 2, previous methods often separate the identity (ID) guidance into a separate attention layer, where the result of the ID attention is added to the output of the original cross-attention layer guided by the text.
 In contrast, we find that our approach of concatenating the keys and values from the face identity directly into the original cross-attention layer provides stronger and more consistent ID guidance.
@@
 Instead, the concatenated keys and values are seamlessly integrated during the multiplication process, and the hidden state is restored to its original size without any additional projections.
-By concatenating the keys and values, our method ensures that the identity information is effectively integrated into the attention mechanism, leading to more accurate and consistent preservation of facial characteristics in the generated images.
-simplifies the architecture but also improves the overall performance and efficiency of the model.
-3.3 MULTIPLE REFERENCE
-Overall, we believe that a single image clearly showing the identity instance is sufficient for identitypreserving generation.
-However, there are cases where using multiple references can yield better results, especially when the face in the adapter is occluded or different features are displayed across different reference images.
+3.3 Multiple References
+Figure 3: Comparison between using n references to guide the generation by our methods, first column represents the references identities, second column is the guidance using only one image of the references, third is using two references And fourth using all three, we notice that generated face is a mix of the features in the references.
+Overall, we believe that a single image clearly showing the identity instance is sufficient for identity-preserving generation.
+However, there are cases where using multiple references can yield better results, especially when the face in the reference input is occluded or different features are displayed across different reference images.
 Our method can easily accommodate the use of multiple references.
-For identities $f1,\dots,fn$ with corresponding hidden states $h1,\dots,hn$, the keys and values can be obtained using Equation 2.
-Equation 3 is then modified as follows:
-$K`=[K_{c};K_{f1};\dots;K_{fn}],V`=[V_{c};V_{f1};\dots;V_{fn}]$ (6)
+For identities $f_{1},\dots,f_{n}$ with corresponding hidden states $h_{1},\dots,h_{n}$, the keys and values can be obtained using Equation 2 and Equation 3 is then modified as follows:
+$K^{{}^{\prime}}=[K_{c};K_{f_{1}};\dots;K_{f_{n}}],V^{{}^{\prime}}=[V_{c};V_{f_{1}};\dots;V_{f_{n}}]$ (6)
 This modification allows for the integration of any number of identities without the need for further training or fine-tuning.
 We showcase sample results in Figure 3, demonstrating the effectiveness of using three identities for improved generation.
-The same principle can be applied to generate an image with an identity that is a mixture of distinct identities, a process commonly known as face morphing.
-More details on this application can be found in the appendix.
+The same principle can be applied to generate an image with an identity that is a mixture of distinct identities, a process commonly known as face morphing (Fu & Damer, 2022), It is important to note that the order of the identities matters, as the attention mechanism in SD (Rombach et al., 2022) does not account for this.
 By extending our method to handle multiple references, we enhance the flexibility and robustness of the identity-preserving generation process.
@@
 3.4 Multiple Identities
+Figure 4: Multiple face identities image generation using our proposed method, first row represents the three input distinct identities, following row contains sample results for different prompts, depthmap based ControlNet (Zhang et al., 2023) was used for the pose.
 Handling multiple identities in image generation remains an open challenge.
-In this work, we propose a novel solution, that combines our previously mentioned method to address this issue.
-The goal is to generate an image containing (n) distinct identities ${f1,\dots,fn}$ based on input references and a given prompt.
-Given masks ${m1,\dots,mn}$ where mi indicates The location of The face for identity i, We scale each mask and feed to each cross-attention layer In The UNet, The scaling issue done based one the scale of the hidden state input To the attention and for some layer with scales si We have masks ${m1si,\dots,mnsi}$.
-For each of the face identities We can get get keys And values And combined them similar to Equation 6, if We applied attention like this we will obtain faces that are mix off all identities features (i.e.
-face morphing), we instead use a customized attention mask M, where The mask can is defined as:
-$M_{q,k}=\begin{cases}-\inf&\text{if }isFaceKey(k)\times identity(k)=i\wedge\exists m_{j}:m_{j_{q}}=true\wedge j\neq i\\ 1&\text{otherwise }\end{cases}$ (7)
-where q indicates the query and k is the key.
-This approach ensures that all identities are present in the generated image.
-We demonstrate the effectiveness of our method with graphical results showing example input identities in Figure 4.
+In this work, we propose a solution, that combines our previously mentioned method to address this issue.
+The goal is to generate an image containing (n) distinct identities $f_{1},\dots,f_{n}$ based on input references and a given prompt.
+Given masks $M_{1},\dots,M_{n}$ where $M_{i}$ indicates the location mask of the face for identity i, we scale each mask and feed to each cross-attention layer in the UNet, the scaling is done based on the scale of the hidden state input to the attention and for some layer with scales s the reshaped and scaled masks are ${M_{1}}^{(s)},\dots,{M_{n}}^{(s)}$, for each of the face identities, we can get keys and values that are combined following Equation 6, if we applied attention like this we will obtain faces that are mix of all identities features (i.e.
+face morphing), we instead use a customized attention mask, denoted as ${\bm{A}}$, where the mask is defined as:
+${\bm{A}}_{q,k}=\begin{cases}-\inf&\text{if }isFaceKey(k)\wedge identity(k)=i\wedge\exists{M_{j}}^{(s)}:{M_{j_{q}}}^{(s)}=true\wedge j\neq i\\ 1&\text{otherwise }\end{cases}$ (7)
+where q indicates the query and k is the key, this restricts the guidance for each face reference to a single, specific face, it also ensures that all identities are present in the generated image.
+We demonstrate the effectiveness of our method with graphical results using distinct input identities in Figure 4.
 These results highlight the success of our approach in generating images conditioned on multiple face identities.
-Figure 3: The overall pipeline of our proposed method.
-We first concatenate the target face image with the image to be generated, We modify the cross attention layers to include new keys any values layers their output is concatenated with the output of the original ones from text prompt.
 4 Experiments
-Figure 4: Multiple face identities image generation using our proposed method, first row represents the three input distinct identities, following row contains sample results for different prompts, depthmap based ControlNet Zhang et al.
-(2023) was used for the pose.
-Figure 5: Comparison with other SOTA methods, with prompt used as ”astronaut in a garden” and the first row represents the input target faces identities, the following rows are results from IP-faceAdapter Ye et al.
-(2023), IP-face-Adapter-plus Ye et al.
-(2023) and ours respectively.
-Identities are selected from ffhq datasetKarras et al.
-(2019)
+Figure 5: Comparison with other SOTA methods, with prompt used as “astronaut in a garden” and the first row represents the input target faces identities, the following rows are results from IP-faceAdapter (Ye et al., 2023), IP-face-Adapter-plus (Ye et al., 2023) and ours respectively.
+Identities are selected from ffhq dataset (Karras et al., 2019).
 4.1 Experiment Settings
-In our work, we utilize Stable Diffusion Rombach et al.
-(2022) as the base model, specifically the Realistic Vision v4.0 checkpoint1 , we train a subset from the LAION-face dataset Zheng et al.
-(2022) consisting of approximately 80000 text-image pairs.
-Our training and all the experiments are done on 2 Tesla V100 GPUs (30 GB) with batch size of 1 per GPU.
-For the other models compared in our work, we use InstantID Wang et al.
-(2024) which uses SDXL model (offering an advantage over our method), as for IPA-FaceId, we use IPA-FaceID-plusv2 Ye et al.
-(2023) and both using Realistic Vision v4.0 checkpoint as the base model, for judging the CLIP score, we generate listing of 100 identity conditioned on different prompts using LLMs.
-We provide two versions of our method, the second version is trained for more timestep and larger dataset.
+In our work, we utilize Stable Diffusion (Rombach et al., 2022) as the base model, specifically the Realistic Vision v4.0 checkpoint [^1] https://huggingface.co/SG161222/Realistic_Vision_V4.0_noVAE.
+We train on a subset of the LAION-face dataset (Zheng et al., 2022), consisting of approximately 80,000 text-image pairs.
+All training and experiments are conducted on two Tesla V100 GPUs (32 GB), with a batch size of 1 per GPU.
+For comparison with other models, we use InstantID (Wang et al., 2024), which employs the SDXL model (providing an advantage over our method).
+For IPA-FaceId, we use IPA-FaceID-plusv2 (Ye et al., 2023), both utilizing the Realistic Vision v4.0 checkpoint as the base model.
+To evaluate the CLIP score, we generate a list of 100 identities conditioned on different prompts using LLMs.
+We provide two versions of our method, with the second version trained for more timesteps and on a larger dataset.
 4.2 Evaluation Metrics
-Previous research on identity-preserving imagine generation tasks Wang et al.
-(2024); Ye et al.
-(2023); Papantoniou et al.
-(2024) lacks a defined metric for assessing identity preservation.
+Previous research on identity-preserving imagine generation tasks (Wang et al., 2024; Ye et al., 2023) lacks a defined metric for assessing identity preservation.
 We employ cosine distance to assess the ID-conditioning capability of various approaches in our study.
@@
 We compare the face embedding of the reference image with those from the created image; a reduced distance signifies superior ID-consistent generating performance.
-Additionally, we calculate the CLIP score using the CLIP modelRadford et al.
-(2021) to assess content alignment in face image generation across several scenarios with varied text prompts.
-The higher CLIP score indicates high  [^1] https://huggingface.co/SG161222 text-image alignment.
-For the evaluation of image quality, we employ SSIM and PSNR to compare the generated images with the reference images.
-In order to measure the face fidelity, we focus on calculate those metrics based on the face region only by cropping the faces from generated images.
-Figure 6: Comparison with other SOTA methods, the first column are the input identities along with input styles, From left to right, we show the results of IPA Ye et al.
-(2023), IPA-FaceID-SDXL, IPAFaceID-Plus, InstantId and our’s, it’s noticed that previous methods unlike our’s can suffer from lack of Face fidelity or style degradation.
+Additionally, we calculate the CLIP score using the CLIP model (Radford et al., 2021) to assess content alignment in face image generation across several scenarios with varied text prompts, the higher CLIP score indicates high text-image alignment.
+For the evaluation of image face quality, we employ SSIM and PSNR to compare the generated images with the reference images.
+In order to measure the face fidelity, we focus on calculating those metrics based on the face region only by cropping the faces from the generated images.
+Figure 6: Visual comparison with other SOTA methods in different generation styles.
+The first column are the input identities along with input styles, from left to right, we show the results of IPA (Ye et al., 2023), IPA-FaceID, IPA-FaceID-SDXL, IPA-FaceID-Plus, InstantID (Wang et al., 2024) and ours.
+It’s noticed that previous methods suffer from the lack of face fidelity or style degradation.
+Images are selected from InstantID (Wang et al., 2024).
 4.3 Qualitative Results
-In this section, we mainly evaluate our proposed method for visual comparison in terms of ID preservation.
-As showns in Figure 5, we generate images based on multiple identity face images with the same text prompt ”astronaut in a garden”.
-IPA-FaceID tend to generate face images with low ID preservation and it sometimes fails to maintain the gender of the reference face image.
-On the other hand, IPA-FaceID-Plus has better performance on maintaining the gender, age of input identity, however, it still loses some original facial attributes such as facial emotion and hairstyles.
-In the contrast, our method generates face images have high level of ID preservation (preserving facial expression, gender, age and detailed feature such as hairstyles and skincolor).
-Specifically, in order to investigate the accurate facial expression generation, we generate images with different text prompts like ”Anger”, ”Cry”, ”Laugh” to control the face emotion of generation.
-As depicted in Figure 1, InstantID generates images with less emotion alignment with associated text prompt while IPA-FaceID-Plus can generate intend emotion.
-However, IPA-FaceID-Plus has less control of ID preservation with exaggerated face expression in random styles (noticeable changes in facial structures and diverse backgrounds).
-Our method provides the strong control of facial expression based on corresponding text prompt.
-Especially in the case of ”Cry” prompt, even though the reference image already with ‘smile’ or ‘laugh’ facial expression, our method still can address the importance of text prompt to the final image generation.
-Our method produces ID-consistent face images while has more natural-looking face expression representations without suppressing the contribution of text prompt to the generated images.
+In this section, we evaluate our proposed method for visual comparison in terms of identity (ID) preservation.
+As shown in Figure 5, we generate images based on multiple identity face images with the same text prompt “astronaut in a garden”.
+IPA-FaceID tends to generate face images with low ID preservation and sometimes fails to maintain the gender and age of the reference face image.
+On the other hand, IPA-FaceID-Plus performs better in maintaining the gender and age of the input identity; however, it still loses some original facial attributes, such as skin tone and hairstyles.
+In contrast, our method generates face images with a high level of ID preservation, preserving facial expressions, gender, age, and detailed features such as hairstyles and skin color.
+Specifically, to investigate accurate facial expression generation, we generate images using different text prompt styles like “Anger”, “Cry”, and “Laugh” to control the facial emotion in the generated images.
+As depicted in Figure 6, InstantID generates images with less emotional alignment to the associated text prompts, while IPA-FaceID-Plus can generate the intended emotion.
+However, IPA-FaceID-Plus has less control over ID preservation, producing exaggerated facial expressions in random styles.
+Our method provides strong control of facial expressions based on the corresponding text prompt, especially in the case of the “Cry” prompt, even if the reference image already has a ‘smile’ or ‘laugh’ facial expression, our method still emphasizes the importance of the text prompt in the final image generation.
+Our method produces ID-consistent face images while delivering more natural-looking facial expression representations, without suppressing the contribution of the text prompt to the generated images.
 In Figure 6, we aim to compare the ID preservation in different image styles generation.
-Same as previous work Wang et al.
-(2024), we use a set of text prompts to indicate various image styles like ”Line art”, ”Faded Film” and ”Watercolor”.
+Same as previous work InstantID (Wang et al., 2024), we use a set of text prompts to indicate various prompt styles like “Line art”, “Faded Film” and “Watercolor”.
 Our method is able to generate comparable results with other SOTA methods while maintaining high degree of face and style alignment together.
+Table 1: Quantative evaluations of ID preservation.
+Cosine distance $\downarrow$
+Method VGG-faces Facenet512 ArcFace
+IPA-FaceID 0.60 0.27 0.55
+IPA-FaceID-Plus 0.36 0.25 0.39
+InstantID 0.42 0.27 0.38
+$Ours_{v1}$ 0.33 0.22 0.34
+$Ours_{v2}$ 0.25 0.18 0.28
 4.4 Quantitative Results
-Besides visual comparisons, we also numerically evaluate our method based on CLIP score, SSIM, PSNR and cosine distance using three models namely VGG-Face, Facenet512 and ArcFace.
-We utilize a subset of FFHQ Karras et al.
-(2019) dataset high quality face images as the input identities in these experiments.
+Besides visual comparisons, we also numerically evaluate our method based on CLIP score, SSIM, PSNR and cosine distance using deepface (Serengil & Ozpinar, 2024; 2020) utilizing three models namely VGG-Face, Facenet512 and ArcFace.
+We utilize a subset of FFHQ (Karras et al., 2019) dataset high quality face images as the input identities in these experiments.
 Table 1 shows the numerical results of cosine distance on different methods.
 Firstly, in terms of the face alignment, it’s clear that our methods achieve significantly higher alignment than the other methods on all face recognition models.
-The faces in our generated image maintain high level of face features from original reference identity and can be easily used for accessing different face recognition systems.
-In Table 2, we judge The face quality And prompt alignment offering different methods, IPA-FaceId has the highest CLIP score, however it’s ID similarity is very low compared together ours and other methods, Ours v1 And IPA-FaceID-Plus maintains almost similar CLIP score with IPA-FaceID-Plus being slightly better, Ours v2 is worse In terms of CLIP but it achieves significantly higher identifier similarity.
-We exclude InstantId Wang et al.
-(2024) from CLip Score due to the fact that it requires pose conditioning And the base model for the public checkpoint issue SDXL based, which will make The comparison not plausible.
-We also compare The face quality using PSNR and SSIM metrics, our method achieves benefits result on both further demonstrating that our method is capable of generating high quality realistic faces with high ID fidelity.
+The faces in our generated image maintain high level of face features from original reference identity and can be easily used for deceiving different face recognition systems.
+In Table 2, we judge the face quality and prompt alignment offering different methods, IPA-FaceId has the highest CLIP score, however its’ ID similarity is very low compared to ours and other methods, ours v1 and IPA-FaceID-Plus maintains almost similar CLIP score with IPA-FaceID-Plus being slightly better, ours v2 is worse in terms of CLIP, but it achieves significantly higher identity similarity.
+We exclude InstantID (Wang et al., 2024) from CLIP score due to the fact that it requires pose conditioning and the base model for the public checkpoint is SDXL based, which will make the comparison not plausible.
+We also compare the face quality using PSNR and SSIM metrics, our method achieves best result on both, further demonstrating that our method is capable of generating high quality realistic faces with high ID fidelity.
 5 limitations and future work
 We have observed that fine facial features tend to degrade when the face occupies a small portion of the overall image.
-This issue primarily stems from the limitations of Stable Diffusion Rombach et al.
-(2022).
-Additionally, when our method is used in conjunction with ControlNet Zhang et al.
-(2023), it results in reduced identity fidelity, which can begin resolved by finetuning.
+This issue primarily stems from the limitations of Stable Diffusion (Rombach et al., 2022).
+Additionally, when our method is used in conjunction with ControlNet (Zhang et al., 2023), it results in slight reduction in identity fidelity, which can be resolved by light fine-tuning or reducing the conditioning strength.
 In our future work, we aim to address these challenges by enhancing the ability to maintain fine details and improving control in scenarios involving the generation of multiple identities.
@@
 Our approach demonstrates robust generalization across various facial views and expressions.
-Additionally, we propose innovative solutions for handling multiple references and multi-identity generation, effectively bridging encoder domain gaps and leveraging multi-scale guidance for enhanced performance.
+Additionally, we propose innovative solutions for handling multiple references and multi-identity generation, effectively bridging encoder domain gaps and leveraging multi-reference guidance for enhanced performance.
 Our method achieves new state-of-the-art results in preserving identity similarity, as evidenced by our extensive experimental evaluations.
 Furthermore, it maintains good alignment with given prompts, setting a new benchmark in the field of identity-preserving image generation.
-Our proposed methods is capable of high prompt alignment with multi-views Face deduction from one sample input, while preserving the target identity with high fidelity 2, we can clearly see that other methods can fail in identity preservation on certain difficult reference inputs and loses prompt alignment, generating same face pose And landmarks is also another issue, we can see that our proposed pipeline and method of training overcome these issue as shown in 5.
+Table 2: Quantative evaluations of face quality and prompt alignment.
+Method CLIP Score $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
+IPA-FaceID 25.23 27.94 0.36
+IPA-FaceID-Plus 23.76 28.04 0.44
+InstantID - 27.93 0.33
+$Ours_{v1}$ 23.28 28.25 0.41
+$Ours_{v2}$ 22.51 28.40 0.47